sonia had a reccuring deposit account in a bank and deposited Rs.600 per month for 2and ½ years.if the rate of interest waas 10% per annum.,find the maturity value of this account.

Grade:10

1 Answers

Raj ojha
5 years ago
Given:p=Rs.600r=10%n=5/2years =30monthsmv=p*n+p*n(n+1)/2*12*r/100mv=600*30+600*30(30+1)/2*12*10/100mv=18000+18000*31/24*10mv=18000+2325mv=Rs.20325